The states of Hawaii and Pennsylvania spend a total of $60.9 million for tourism. The state of Hawaii spends 8.1 million less than twice the amount of money that Pennsylvania spends.
Find the amount that each state spends on tourism

Answer by checkley75(3666)

LET X BE THE SPENDING FOR PENN.
THEN THE SPENDING FOR HAWAII=2X-8.1
TOGETHER THEY SPEND 60.9.
THUS WE HAVE X+(2X-8.1)=60.9
X+2X-8.1=60.9
3X=60.9+8.1
3X=69
X=69/3
X=$23 MILLION IS SPENT BY PENN.
2*23-8.1=46-8.1= 37.9 MILLION IS SPENT BY HAWAII.
PROOF
23+37.9=60.9
60.9=60.9